# Convergence and quantum advantage of Trotterized MERA for strongly-correlated systems

### v2

@@ 1 Introduction
Due to the quantum advantage [1, 2] for certain tasks, quantum computers are expected to lead to a revolution in information technology. While there has been enormous theoretical and experimental progress, quantum computation currently still faces a number of challenges, including the limited number of qubits, inevitable noise, and insufficient fidelity [3, 4]. The simulation of quantum matter is less demanding and, hence, a promising field of application for the current era of noisy intermediate-scale quantum (NISQ) devices [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Compared with classical simulation techniques, quantum simulation methods can have substantially lower time complexity and may, hence, lead to new insights on quantum many-body physics in the near future.
@@
In this paper, we try to establish a quantum advantage of the TMERA VQE over classical MERA algorithms and describe further improvements for its practical implementation. Sec. 4 describes and analyzes different schemes for initializing the TMERA optimization and driving it to convergence with the goal of avoiding local minima. Iteratively increasing the number of variational parameters can substantially improve the performance. One can also start from system parameters with a well-known low-entangled ground state and, during the optimization, then move on certain paths through the space of model parameters to obtain the actual ground state of interest. In the central Secs. 5 and 6, we analyze the quantum-computational complexity of TMERA VQE and compare to the classical computation costs of (unconstrained) full MERA (fMERA) as well as TMERA methods, taking into account the stochastic nature of quantum measurements and considering the classical methods employing exact energy gradients (EEG) [44] as well as a variational Monte Carlo (VMC) scheme [45, 46]. For several critical one-dimensional spin-$s$ quantum magnets, we find a polynomial quantum advantage for the TMERA VQE over the classical MERA methods. In the considered examples, the quantum advantage increases with increasing spin quantum number $s$. We determine algorithmic phase diagrams which suggest considerably larger quantum advantages for systems in $D\geq 2$ spatial dimensions. Small two-qubit rotation angles are desirable for implementations on present-day quantum devices. In Sec. 7, we add an angle penalty term to the energy functional and find that the average angle can be reduced considerably without substantially affecting energy accuracies. The initial TMERA proposal [10] chose all TMERA tensors as brick-wall circuits (a.k.a. alternating layered ansatz) with generic two-qubit gates acting on nearest neighbors. In Sec. 8, we replace these brick-wall circuits by parallel random-pair circuits (PRPC). While PRPC contain long-range two-qubit gates, they do not seem to outperform brick-wall circuits, as long as bond dimensions are relatively small. Section 9 summarizes the findings and comments on the broader context.
@@ 5 Computational time complexity
where $U_{\tau,k}$ denotes one of the MERA tensors in layer $\tau$. The computation time follows a power law $\mathcal{O}(\chi^{r})$ with rather large exponents $r$ especially in $D\geq 2$ spatial dimensions. This holds regardless of whether the optimization is based on Riemannian gradients [97, 10] or the traditional fixed-point-iteration SVD scheme [44]. The costs for EEG-based TMERA optimizations depend additionally on the number of Trotter steps $t$ needed to achieve a certain energy accuracy (6). In the VMC-based fMERA and TMERA optimizations, energy gradients (8) are not determined exactly but sampled stochastically, where, very similarly to the VQE scheme, one propagates pure causal-cone states in the MERA preparation direction and applies projective measurements to the sites that leave the causal cone in the layer transitions [45, 46]. Hence, the associated cost attains an additional factor $1/\epsilon^{2}$ corresponding to the required number of samples. As will also be confirmed in the benchmark simulations, for critical models, the energy accuracy and optimal number of Trotter steps scale algebraically with respect to the bond dimension $\chi$,
@@ 6 Quantum advantage
Algorithmic phase diagrams. — Based on the scaling of computation costs in Eq. (11) and Table 1, Fig. 5 provides algorithmic phase diagrams. As a function of the scaling exponents $\beta$ and $p\leq p_{\text{max}}$ for the energy accuracy and number of Trotter steps, the diagrams show which of the five considered methods has the lowest computation costs per optimization step in the large-$\chi$ regime. For all six types of MERA structures (specified in Sec. 5), the TMERA VQE dominates the realistic $(\beta,p)$ ranges. For the 1D modified binary MERA (Fig. 5b), in particular, all critical groundstate problems from Sec. 3 lie in the region where TMERA VQE is most efficient. The spin-1/2 XXZ chain in the critical XY phase ($|\Delta|\leq 1$) happens to fall directly onto the transition line from the TMERA-VQE region to a region where fMERA and TMERA EEG have the same and best cost scaling as already suggested by the results in Fig. 3. Generally, the classical algorithms only prevail in relatively small or unrealistic $(\beta,p)$ regions with fMERA VMC at large $p$ and small $\beta$ as well as fMERA and/or TMERA EEG at large $p$ and large $\beta$. Note, for example that, in the diagram for the 2D quaternary MERA [105], no classical algorithm can compete with the VQE for $0\leq\beta\leq 10$ (see Fig. 5d). Lastly, according to this analysis, the classical TMERA VMC method [46] is not competitive for any of the considered 1D and 2D MERA.
We can conclude that the observed polynomial advantage of TMERA VQE over the classical algorithms is predominantly due to replacement of the costly classical tensor contractions with the execution of Trotter quantum circuits on the quantum computer. Neither the tensor Trotterization (TMERA) nor importance sampling (VMC), when applied separately or in combination on classical computers, can outperform the VQE for realistic values of the energy and Trotter scaling exponents. The separation is more pronounced for systems in $D\geq 2$ dimensions due to the rapid increase of MERA contraction costs for higher-dimensional systems. In contrast to the classical EEG methods, quantum computers produce probabilistic measurement outcomes. However, the results show that drawbacks of the resulting necessity to sample energy gradients are outweighed by the avoidance of high classical tensor contraction costs.
@@ 7 Reducing rotation angle amplitudes
Figure 6: Effect of large-angle penalties. The plot shows results for the spin-1/2 XXX model using homogeneous modified binary TMERA with $T=6$, $t=8$, and $q=3$. The modified objective function (14) comprises the energy density and a term penalizing large rotation angles. The three lower curves show the maximum, mean value, and minimum of energy density accuracy for a few hundred randomly initialized TMERA optimizations. The displayed average two-qubit rotation angle is obtained from the converged lowest-energy TMERA.
@@ 8 Random circuit TMERA tensors
So far, we chose the TMERA tensors as brick-wall circuits with Trotter gates acting on nearest-neighbor qubits. The expressiveness and efficiency of the ansatz might be improved for alternative choices. In ion-trap systems one could, for example, use the all-to-all connectivity [109, 110] and work, instead, with parallel random-pair circuits (PRPC) as illustrated in Fig. 1c. Each Trotter step consists again of two Trotter-gate coverings. In each covering, $n/2$ two-qubit gates act on disjoint qubit pairs, randomly chosen from the $n$-qubit support of the TMERA tensor. For the spin-1/2 Heisenberg antiferromagnet, Fig. 7 shows a comparison of TMERA accuracies for brick-wall circuit tensors and PRPC tensors. Even though a PRPC contains long-range two-qubit gates and hence allows a faster scrambling, the scaling of the attained accuracies is very similar, at least for this model and relatively small bond dimensions.
@@ 9 Discussion
    Arguments by Kim and Swingle [15] also apply to TMERA and suggest its resilience against noise. This is due to the fact that MERA layer transition channels are generally gapped such that noise does not proliferate but gets damped when transitioning to the next layer.
@@
As an example concerning the latter point, a 1D binary TMERA VQE with bond dimension $\chi=8$ ($q=3$), $T=6$ layers, and $t=2$ Trotter steps requires only 12 qubits and circuit depth 120 when optimizing for low qubit number, or 18 qubits and circuit depth 48 when optimizing for low depth. Also, a 2D $2\times 2\mapsto 1$ TMERA VQE with bond dimension $\chi=4$ ($q=2$), $T=6$ layers, and $t=2$ Trotter steps requires only 28 qubits and circuit depth 624 when optimizing for low qubit number, or 72 qubits and circuit depth 96 when optimizing for low depth.

### v3

@@ 1 Introduction
Due to the quantum advantage [1, 2] for certain tasks, quantum computers are expected to lead to a revolution in information technology. While there has been enormous theoretical and experimental progress, quantum computation currently still faces a number of challenges, including the limited number of qubits, inevitable noise, and insufficient fidelity [3, 4]. The simulation of quantum matter is less demanding and, hence, a promising field of application for the current era of noisy intermediate-scale quantum (NISQ) devices [5, 6, 7, 8, 9, 10, 11, 12, 13, 14]. Compared to classical techniques, quantum simulation methods can have substantially lower time complexity and may, hence, lead to new insights on quantum many-body physics in the near future.
@@
In this paper, we try to establish a quantum advantage of the TMERA VQE over classical MERA algorithms and describe further improvements for its practical implementation. Sec. 4 describes and analyzes different schemes for initializing the TMERA optimization and driving it to convergence with the goal of avoiding local minima. Iteratively increasing the number of variational parameters can substantially improve the performance. One can also start from system parameters with a well-known low-entangled ground state and, during the optimization, then move on certain paths through the space of model parameters to obtain the actual ground state of interest. In the central Secs. 5 and 6, we analyze the quantum-computational complexity of TMERA VQE and compare to the classical computation costs of (unconstrained) full MERA (fMERA) as well as TMERA methods, taking into account the stochastic nature of quantum measurements and considering the classical methods employing exact energy gradients (EEG) [44] as well as a variational Monte Carlo (VMC) scheme [45, 46]. For several critical one-dimensional spin-$s$ quantum magnets, we find a polynomial quantum advantage for the TMERA VQE over the classical MERA methods. In the considered examples, the quantum advantage increases with increasing spin quantum number $s$. We determine algorithmic phase diagrams that suggest considerably larger quantum advantages for systems in $D\geq 2$ spatial dimensions. Small two-qubit rotation angles are desirable for implementations on present-day quantum devices. In Sec. 7, we add an angle penalty term to the energy functional and find that the average angle can be reduced considerably without substantially affecting energy accuracies. The initial TMERA proposal [10] chose all TMERA tensors as brick-wall circuits (a.k.a. alternating layered ansatz) with generic two-qubit gates acting on nearest neighbors. In Sec. 8, we replace these brick-wall circuits by parallel random-pair circuits (PRPC). While PRPC contain long-range two-qubit gates, they do not seem to outperform brick-wall circuits, as long as bond dimensions are relatively small. Section 9 summarizes the findings and comments on the broader context.
@@ 5 Computational time complexity
where $U_{\tau,k}$ denotes one of the MERA tensors in layer $\tau$. The computation time follows a power law $\mathcal{O}(\chi^{r})$ with rather large exponents $r$ especially in $D\geq 2$ spatial dimensions. This holds regardless of whether the optimization is based on Riemannian gradients [97, 10] or the traditional fixed-point-iteration SVD scheme [44]. The costs for EEG-based TMERA optimizations depend additionally on the number of Trotter steps $t$ needed to achieve a certain energy accuracy (6). In the VMC-based fMERA and TMERA methods, energy gradients (8) are not determined exactly but sampled stochastically, where, in analogy to the VQE scheme, one propagates pure causal-cone states in the MERA preparation direction and applies projective measurements to the sites that leave the causal cone in the layer transitions [45, 46]. Hence, the associated cost attains an additional factor $1/\epsilon^{2}$ corresponding to the required number of samples. As will also be confirmed in the benchmark simulations, for critical models, the energy accuracy and optimal number of Trotter steps scale algebraically with respect to the bond dimension $\chi$,
@@ 6 Quantum advantage
Algorithmic phase diagrams. — Based on the scaling of computation costs in Eq. (11) and Table 1, Fig. 5 provides algorithmic phase diagrams. As a function of the scaling exponents $\beta$ and $p\leq p_{\text{max}}$ for the energy accuracy and number of Trotter steps, the diagrams show which of the five considered methods has the lowest computation costs per optimization step in the large-$\chi$ regime. For all six types of MERA structures (specified in Sec. 5), the TMERA VQE dominates the realistic $(\beta,p)$ ranges. For the 1D modified binary MERA (Fig. 5b), in particular, all critical groundstate problems from Sec. 3 lie in the region where TMERA VQE is most efficient. The spin-1/2 XXZ chain in the critical XY phase ($|\Delta|\leq 1$) happens to fall directly onto the transition line from the TMERA-VQE region to a region where fMERA and TMERA EEG have the same and best cost scaling as already suggested by the results in Fig. 3. Generally, the classical algorithms only prevail in relatively small or unrealistic $(\beta,p)$ regions with fMERA VMC at large $p$ and small $\beta$ as well as fMERA and/or TMERA EEG at large $p$ and large $\beta$. Note for example that, in the diagram for the 2D quaternary MERA [105], no classical algorithm can compete with the VQE for $0\leq\beta\leq 10$ (see Fig. 5d). Lastly, according to this analysis, the classical TMERA VMC method [46] is not competitive for any of the considered 1D and 2D MERA.
We can conclude that the observed polynomial advantage of TMERA VQE over the classical algorithms is predominantly due to the replacement of the costly classical tensor contractions with the execution of Trotter quantum circuits on the quantum computer. Neither the tensor Trotterization (TMERA) nor importance sampling (VMC), when applied separately or in combination on classical computers, can outperform the VQE for realistic values of the energy and Trotter scaling exponents. The separation is more pronounced for systems in $D\geq 2$ dimensions due to the rapid increase of MERA contraction costs for higher-dimensional systems. In contrast to the classical EEG methods, quantum computers produce probabilistic measurement outcomes. However, the results show that drawbacks of the resulting necessity to sample energy gradients are outweighed by the avoidance of high classical tensor contraction costs.
@@ 7 Reducing rotation angle amplitudes
Figure 6: Effect of large-angle penalties. The plot shows results for the spin-1/2 XXX model using homogeneous modified binary TMERA with $T=6$, $t=8$, and $q=3$. The modified objective function (14) comprises the energy density and a term penalizing large rotation angles. The three lower curves show the maximum, mean value, and minimum of the energy density accuracy for a few hundred randomly initialized TMERA optimizations. The displayed average two-qubit rotation angle is obtained from the converged lowest-energy TMERA.
@@ 8 Random circuit TMERA tensors
So far, we chose the TMERA tensors as brick-wall circuits with Trotter gates acting on nearest-neighbor qubits. The expressiveness and efficiency of the ansatz might be improved for alternative choices. In ion-trap systems one could, for example, use the all-to-all connectivity [109, 110] and work, instead, with parallel random-pair circuits (PRPC) as illustrated in Fig. 1c. Each Trotter step consists again of two Trotter-gate coverings. In each covering, $n/2$ two-qubit gates act on disjoint qubit pairs, randomly chosen from the $n$-qubit support of the TMERA tensor. For the spin-1/2 Heisenberg antiferromagnet, Fig. 7 shows a comparison of TMERA accuracies for brick-wall circuit tensors and PRPC tensors. Even though a PRPC contains long-range two-qubit gates and hence allows for faster scrambling, the scaling of the attained accuracies is very similar, at least for this model and relatively small bond dimensions.
@@ 9 Discussion
    Arguments by Kim and Swingle [15] also apply to TMERA and suggest their resilience against noise. This is due to the fact that MERA layer transition channels are generally gapped such that noise does not proliferate but gets damped when transitioning to the next layer.
@@
As an example concerning the latter point, a 1D binary TMERA VQE with bond dimension $\chi=8$ ($q=3$), $T=6$ layers, and $t=2$ Trotter steps requires only 12 qubits and circuit depth 120 when optimizing for low qubit number, or 18 qubits and circuit depth 48 when optimizing for low depth. Also, a 2D $2\times 2\mapsto 1$ TMERA VQE with bond dimension $\chi=4$ ($q=2$), $T=6$ layers, and $t=2$ Trotter steps requires only 28 qubits and circuit depth 624 when optimizing for low qubit number, or 72 qubits and circuit depth 96 when optimizing for low depth. First experimental implementations of MERA on ion-trap quantum computers have demonstrated critical groundstate correlation functions [125] as well as a quantum phase transition and the associated change from area-law to log-area-law entanglement scaling [126].